OpenStudy (anonymous):

Help please! :) Racing cyclists use the function r(x) = 4√x to determine the maximum rate, r(x), in miles per hour, to turn a corner of radius x, in feet, without tipping over. What is the maximum rate a cyclist should travel around a corner of radius 18 feet without tipping over? Leave the solution in simplified radical form.

hero (hero):

Just replace x = 18 into the formula for r(x), then evaluate the right side

hero (hero):

r(x) = 4√x r(18) = 4√18 = 4√(9*2) = 4*3√2 = 12√2 miles per hour

OpenStudy (anonymous):

very thankful :) you rock!

hero (hero):

I'm guessing you're just looking for answers, right?

OpenStudy (anonymous):

I do need to show work- but you did that already ;)

hero (hero):

What I mean is, are you interested in understanding the concepts behind these problems?

OpenStudy (anonymous):

Well, I thought that was right, by inputting 18 but I didn't think it was THAT simple and yes, I DO want to learn.